(sin x-1)(tan x+sec x) = cos x
Answer by fractalier(6550)

From
(sin x-1)(tan x+sec x) = cos x
we will FOIL out the left side and get
sin x tan x + sin x sec x - tan x - sec x =
Now sin x sec x = tan x, so
sin x tan x + tan x - tan x - sec x =
sin x tan x - sec x =
Now tan x = sin x / cos x and sec x = 1/cos x, so we have
sin^2 x / cos - 1 / cos x =
(sin^2 x - 1) / cos x =
-cos^2 x / cos x =
- cos x
The right hand side should be -cos x...I think you copied it wrong...